# Modeling local predictive ability using power-transformed Gaussian processes

### v1

@@ Abstract
A Gaussian process is proposed as a model for the posterior distribution of the local predictive ability of a model or expert, conditional on a vector of covariates, from historical predictions in the form of log predictive scores. Assuming Gaussian expert predictions and a Gaussian data generating process, a linear transformation of the predictive score follows a noncentral chi-squared distribution with one degree of freedom. Motivated by this we develop a non-central chi-squared Gaussian process regression to flexibly model local predictive ability, with the posterior distribution of the latent GP function and kernel hyperparameters sampled by Hamiltonian Monte Carlo. We show that a cube-root transformation of the log scores is approximately Gaussian with homoscedastic variance, which makes it possible to estimate the model much faster by marginalizing the latent GP function analytically. Linear pools based on learned local predictive ability are applied to predict daily bike usage in Washington DC.
@@ 1 Introduction
The out-of-sample predictive accuracy of an expert, which could be a formal model or an opinionated human, is typically treated as a fixed quantity to be estimated (Gelman et al.,, 2014). Our aim here is to model local predictive ability, i.e. predictive ability as a function of a vector of covariates $\mathbf{z}$, to capture how the predictive abilities of a set of experts vary over the space of $\mathbf{z}$ (Yao et al.,, 2021; Li et al.,, 2021; Oelrich et al.,, 2021). We aim to obtain the posterior distribution of the local predictive ability to properly quantify the uncertainty of the expert’s predictive ability.
@@
Since measures of predictive ability of experts can be used to combine, or pool, expert predictions (Bates and Granger,, 1969; Hall and Mitchell,, 2007; Geweke and Amisano,, 2011; Yao et al.,, 2018), we call the space spanned by this vector the pooling space (Oelrich et al.,, 2021) and the variables in $\mathbf{z}$ pooling variables. The choice of pooling variables is determined by the decision maker (Lindley et al.,, 1979) making the aggregated prediction, based on what she believes are important determinants of the experts’ local predictive abilities.
@@
Our focus is on the accuracy of an expert’s forecast density (Gelman et al.,, 2014). Density forecasts are typically part of the output from probabilistic models, but are also increasingly reported from human experts. For example, the histogram densities for key macroeconomic variables from the Survey of Professional Forecasters conducted by the Federal Reserve Bank of Philadelphia for the US (Croushore et al.,, 2019) and the European Central Bank for the Euro Area (de Vincent-Humphreys et al.,, 2019) or probabilistic forecasts of temperature and precipitation in meteorology from a combination of subjective experts, historical data and complex deterministic scientific models (AMS Council,, 2008).
@@
A theoretically motivated measure of predictive ability is the expected log predictive density (ELPD), which can be unbiasedly estimated by the average log predictive density evaluations (log scores) on a test dataset. The distribution of the log scores is typically complex however (Sivula et al., 2022b, ), making inferences for ELPD challenging. The likelihood of the observed log scores depends not only on the model/expert but also on the data-generating process (DGP). For example, using a simple model with additive noise will typically work poorly, as when the predictive distribution and the DGP are normal, the log scores will follow a scaled and translated noncentral $\chi^{2}_{1}$ distribution (Sivula et al., 2022a, ).
@@
Further, specifying a model for the relationship between all the covariates and predictive ability is a challenging problem in itself. Certain variables may be expected to be important for predictive ability but the exact functional form to use to model this relationship may be harder to pinpoint. In a situation of little prior information it makes sense for the decision maker to go with a non-parametric model that can capture variations in predictive ability over an arbitrary pooling space in a flexible manner. Oelrich et al., (2021) propose the non-parametric caliper estimator of local predictive ability, which works by averaging log scores locally in pooling space. The caliper estimator is flexible, but has the potential drawback that the inferred local ability surface is discontinuous, and the lack of smoothness may result in large variability for the local estimates.
@@
We here propose a flexible noncentral $\chi^{2}$ Gaussian process (GP) regression, where the latent GP surface models the conditional ELPD over the space of pooling variables in $\mathbf{z}$. A Gaussian process explicitly models the correlation properties, but other than that makes minimal assumptions about the specific relationship between the covariates and the ELPD. The smoothness and other hyperparameters in the covariance kernel are assumed unknown and are inferred jointly with the latent GP surface. A Gaussian process model for local ELPD is attractive since it gives a direct quantification of the uncertainty about the underlying ELPD surface for each expert.
@@
The joint posterior distribution of the conditional ELPD and kernel hyperparameters are sampled by Hamiltonian Monte Carlo (HMC) using Stan (Stan Development Team,, 2022). The dimension of the sampled conditional ELPD grows with the sample size however, and inferences are time-consuming even for moderately large datasets. We show empirically that a cube-root transformation of the log scores is approximately Gaussian with homoscedastic variance. This makes it possible to integrate out the latent GP surface analytically and sample the low-dimensional marginal posterior kernel hyperparameters by HMC, which is much faster and scalable for larger datasets. We show that the ELPD surface can then be easily sampled by standard GP methods (Rasmussen and Williams,, 2006, Chapter 2) combined with analytically available third moments of Gaussian processes. In the more general case when either the expert or the data generating process is non-Gaussian we propose to use a general power transformation of the ELPD and to estimate the power transformation using HMC jointly with the kernel hyperparameters.
@@
The paper proceeds as follows. Section 2 defines local predictive ability and describes the non-central $\chi^{2}$ Gaussian process (GP) regression for estimating it, along with the computationally more efficient GP regression on power-transformed log scores. Section 3 illustrates these methods using a simple simulation study. Section 4 consists of an empirical example analyzing local predictive ability for a set of models in a bike sharing dataset. Section 5 concludes.
@@ 2.1 Local predictive ability
The decision maker uses the expected log predictive density for a new data point $\tilde{y}$ from the data-generating process $p_{*}(y)$ to measure the predictive ability of expert $k$
@@
where $\tilde{y}$ is a new (out-of sample) observation, $p_{*}(\cdot)$ is the density of the data-generating process, and $\mathbf{y}_{k}$ is the training data the expert has access to. The variable of interest, $y$, can be univariate or multivariate, but we will for simplicity assume that it is univariate here.
The decision maker further believes that the predictive ability of the experts vary over a set of pooling variables (Oelrich et al.,, 2021). She is therefore not only interested in determining how good the experts are at making predictions in general, but how good they are at making predictions conditional on a vector of pooling variables, $\mathbf{z}$. We call this conditional predictive ability the local predictive ability of the expert at $\mathbf{z}$
@@ 2.2 Distribution of the log scores
The log scores of the experts, $\ell_{ik}$ for $k=1,\ldots,K$ and $i=1,\ldots,n$, are treated as data by the decision maker (Lindley et al.,, 1979) from which she infers how the ELPD varies over the pooling space. The choice of pooling variables is therefore a pure modeling problem from the viewpoint of the decision maker. The decision maker does not need to know if the experts have access to the pooling variables when making their predictions, but this information can of course affect the modeling choices of the decision maker.
The correct model for local predictive ability is determined by the predictive distribution of the expert, which is typically known, and the data-generating process, which is typically unknown. In order to specify a parametric model for the log scores, we therefore need access to the predictive distribution of the expert. Different predictive distribution will lead to different models of local predictive ability.
Consider the practically important scenario where the expert predictive distributions, as well as the data generating process, are all Gaussian. Specifically, let the predictive distribution of expert $k$ for the data point $y_{i}$ be ${y_{i}\mid\mathbf{y}\sim\mathcal{N}(\mu_{ik},\sigma_{ik}^{2})}$, where $\mathbf{y}$ is the training data used by the expert. The log score is then
$\ell_{ik}\equiv\log p_{k}(y_{i}\mid\mathbf{y})=-\frac{1}{2}\log(2\pi\sigma_{ik}^{2})-\frac{1}{2\sigma_{ik}^{2}}(y_{i}-\mu_{ik})^{2}.$ (3)
If we let the data-generating process be given by ${y_{i}\sim\mathcal{N}(\mu_{i*},\sigma_{i*}^{2})}$, we can rewrite Equation (3) as
$\ell_{ik}\equiv\log p_{k}(y_{i})=-\frac{1}{2}\log(2\pi\sigma_{ik}^{2})-\frac{\sigma_{i*}^{2}}{2\sigma_{ik}^{2}}\left(\frac{y_{i}-\mu_{ik}}{\sigma_{i*}}\right)^{2},$ (4)
and since $(y_{i}-\mu_{ik})/\sigma_{i*}\sim\mathcal{N}(\mu_{i*}-\mu_{ik},1)$, it follows that (Sivula et al., 2022a, )
@@
where $X_{ik}\sim\chi^{2}_{1}(\lambda_{ik})$ with non-centrality parameter $\lambda_{ik}=(\mu_{i*}-\mu_{ik})^{2}/\sigma_{i*}^{2}$, $a_{ik}=-\frac{1}{2}\log(2\pi\sigma_{ik}^{2})$ and $b_{ik}=\sigma_{i*}^{2}/(2\sigma_{ik}^{2})$. Hence, the noncentrality parameter $\lambda_{ik}$ depends on the difference between the predictive mean of the expert and the mean of the data-generating process; the shape of the distribution of the log scores therefore varies dramatically with $\mu_{i*}-\mu_{ik}$ and is typically highly non-Gaussian for all $b$, see Figure 1 for an illustration.
@@
where $\chi^{2}_{1}(\lambda,b)$ denotes a scaled non-central $\chi^{2}$ with one degree of freedom, non-centrality parameter $\lambda$ and scale parameter $b$. A useful simplification of the model in (6) is to assume that $b_{ik}=b_{k}$ for all $i=1,\ldots,n$, i.e. that the variance ratio $\sigma_{i*}^{2}/\sigma_{ik}^{2}$ is constant over observations for a given expert.
2.3 Scaled non-central $\chi^{2}$ Gaussian process regression
The setup with Gaussian data generating process and experts in the previous subsection suggests a scaled non-central $\chi^{2}$ regression for modeling the local predictive ability of experts. We will here propose a nonparametric Bayesian model where the log of the noncentrality parameter is modeled by a Gaussian process over the pooling space.
A Gaussian process (GP) generalizes the Gaussian distribution from a random variable to a stochastic process (Rasmussen and Williams,, 2006). In simple terms, a Gaussian process is an infinite sequence of function values with different inputs $\{f(\mathbf{z}_{1}),f(\mathbf{z}_{2}),f(\mathbf{z}_{3}),\dots\}$, any subset of which follows a multivariate Gaussian distribution. Each GP has a covariance function, or kernel, which depends on the input values and specifies how the covariance matrix for each subset of the sequence should be constructed. The specific covariance function we select determines our prior regarding how smoothly the outputs vary over the input space. The covariance function we use in the examples and applications in this paper is called the squared exponential with automatic relevance determination (ARD) (Neal,, 1996)
@@ 2.3 Scaled non-central $\chi^{2}$ Gaussian process regression
as the length scale determines how quickly we assume the function values to vary with regards to a specific input, and the more smoothly the function value varies with regards to a specific input, the more relevant that input can be considered.
Our non-central $\chi^{2}$ Gaussian process model for the log scores is
@@
where $\mathcal{N}^{+}(\cdot,\cdot)$ denotes a normal distribution truncated at 00 Note that $b_{k}>0$, and $b_{k}=1/2$ when the variance of the DGP and the expert match. Further, note that the GP is on the log of the noncentrality parameter, ensuring that $\lambda_{k}(\mathbf{z})$ stays non-negative. Since the mean of a $\chi^{2}_{1}(\lambda,b)$ variable is $b(1+\lambda)$, the Gaussian process in (8) directly implies a flexible prior for the mean function of the log scores, i.e. the local ELPD. We will refer to the model in (8) as $\operatorname{GP}(\chi^{2}_{1})$. The mean function $\mu_{k}(\cdot)$ and the covariance kernel $g_{k}(\cdot,\cdot)$ are subscripted by $k$ since their hyperparameters are allowed to be different across the $K$ experts. Any valid kernel function can be used. Let $\boldsymbol{\theta}_{k}$ be the vector with all the hyperparameters in the mean and kernel function of the GP prior.
The decision maker wants to compute the posterior distribution of the local ELPD $\operatorname{\eta}_{k}(\tilde{\mathbf{z}})=\operatorname{E}_{p_{*}}(\tilde{\ell}_{k})=a_{ik}-b_{k}(1+\lambda_{k}(\tilde{\mathbf{z}}))$ at a new point $\tilde{\mathbf{z}}$ in pooling space for each expert. This is achieved by sampling from the joint posterior
@@
where $\boldsymbol{\ell}^{\prime}_{k}=(\ell^{\prime}_{1k},\ldots,\ell^{\prime}_{nk})^{\top}$ and $\boldsymbol{\lambda}_{k}=(\lambda_{k}(\mathbf{z}_{1}),\ldots,\lambda_{k}(\mathbf{z}_{n}))^{\top}$ is the vector of non-centrality parameters in the training data. We sample the potentially high-dimensional posterior in (9) using Hamiltonian Monte Carlo (HMC) in Stan (Stan Development Team,, 2022; Gabry and Češnovar,, 2020).

### v2

# Modeling local predictive ability with power-transformed Gaussian process regression

@@ Abstract
A Gaussian process is proposed as a model for the posterior distribution of the local predictive ability of a model or expert, conditional on a vector of covariates, from historical predictions in the form of log predictive scores. Assuming Gaussian expert predictions and a Gaussian data generating process, a linear transformation of the predictive score follows a noncentral chi-squared distribution with one degree of freedom. Motivated by this we develop a noncentral chi-squared Gaussian process regression to flexibly model local predictive ability, with the posterior distribution of the latent GP function and kernel hyperparameters sampled by Hamiltonian Monte Carlo. We show that a cube-root transformation of the log scores is approximately Gaussian with homoscedastic variance, making it possible to estimate the model much faster by marginalizing the latent GP function analytically. A multi-output Gaussian process regression is also introduced to model the dependence in predictive ability between experts, both for inference and prediction purposes. Linear pools based on learned local predictive ability are applied to predict daily bike usage in Washington DC.
@@ 1 Introduction
Combining predictions from several models or experts, be they point forecast (Bates and Granger,, 1969) or forecast distributions (Yao et al.,, 2018), has a long history in statistics and econometrics; see Wang et al., (2023) for a recent review. A forecast combination can better quantify total uncertainty (Draper,, 1995) and improve accuracy (Wang et al.,, 2023). Some prominent example of aggregation methods include weighted averages with weights based on past forecast errors (Bates and Granger,, 1969), posterior model probabilities (Hoeting et al.,, 1999) and historical performance of the experts’ density forecasts (Hall and Mitchell,, 2007; Geweke and Amisano,, 2011).
@@
The out-of-sample predictive accuracy of an expert, which could be a formal model or a human, is typically treated as a fixed quantity to be estimated (Gelman et al.,, 2014). Our aim here is to model local predictive ability, i.e. predictive ability as a function of a vector of covariates $\mathbf{z}$, to capture how the predictive abilities of a set of experts vary over the space of $\mathbf{z}$ (Yao et al.,, 2021; Li et al.,, 2023; Oelrich et al.,, 2024). We aim to obtain the posterior distribution of the local predictive ability to properly quantify the uncertainty of the expert’s predictive ability.
@@
Since one use of measures of predictive ability is to combine, or pool, expert predictions, we call the variables in $\mathbf{z}$ pooling variables and the space of $\mathbf{z}$ the pooling space (Oelrich et al.,, 2024). The choice of pooling variables is determined by the decision maker (Lindley et al.,, 1979) making the aggregated prediction, based on what she believes are important determinants of the experts’ local predictive abilities. In an economic application a measure of the business cycle or a survey of expectations could for example be a pooling variable, if the forecasting experts are expected to have different forecasting accuracy depending on where the economy is in the business cycle.
@@
Our focus is on the accuracy of an expert’s forecast density (Gelman et al.,, 2014). Density forecasts are typically part of the output from probabilistic models, but are also increasingly reported from human experts. For example, the histogram densities for key macroeconomic variables from the Survey of Professional Forecasters conducted by the Federal Reserve Bank of Philadelphia for the US (Croushore and Stark,, 2019) and the European Central Bank for the Euro Area (de Vincent-Humphreys et al.,, 2019) or probabilistic forecasts of temperature and precipitation in meteorology from a combination of subjective experts, historical data and complex deterministic scientific models (AMS Council,, 2008).
@@
A theoretically motivated measure of predictive ability is the expected log predictive density (ELPD), which can be unbiasedly estimated by average log predictive density evaluations (log scores) on a test dataset. The distribution of the log scores is typically complex however (Sivula et al., 2022b, ), making inferences for ELPD challenging. The likelihood of the observed log scores depends not only on the model/expert but also on the data-generating process (DGP). For example, using a simple model with additive noise will typically work poorly, as when the predictive distribution and the DGP are normal, the log scores will follow a scaled and translated noncentral $\chi^{2}_{1}$ distribution (Sivula et al., 2022a, ).
@@
Further, specifying the relationship between the set of covariates and predictive ability is a challenging problem in itself. Certain variables may be expected to be important for predictive ability but the exact functional form may be harder to pinpoint. In a situation of little prior information it makes sense for the decision maker to use a non-parametric model that can capture variations in predictive ability over an arbitrary pooling space in a flexible manner. Oelrich et al., (2024) propose the non-parametric caliper estimator of local predictive ability, which averages log scores locally in pooling space. The caliper estimator is flexible, but has the potential drawback that the inferred local ability surface is discontinuous, and the lack of smoothness may result in large variability for the local estimates.
@@
We here propose a flexible noncentral $\chi^{2}$ Gaussian process (GP) regression, where the latent GP surface models the conditional ELPD over the space of pooling variables in $\mathbf{z}$. A Gaussian process explicitly models the correlation properties, but makes otherwise minimal assumptions about the specific relationship between the covariates and the ELPD. The smoothness and other hyperparameters in the covariance kernel are assumed unknown and are inferred jointly with the latent GP surface. A Gaussian process model for local ELPD is attractive since it gives a direct quantification of the uncertainty about the underlying ELPD surface for each expert.
Separate GPs can be estimated for each expert, implicitly assuming independence when forecasts are pooled based on local predictive ability. However, correctly capturing any dependence structure between experts is useful both from a pure inference standpoint, as it allows us to better understand the relationship between the predictions of the experts, as well as for creating better aggregate predictions (McAlinn et al.,, 2020). The latter point is clearly illustrated by Winkler, (1981), who shows that an expert with relatively low predictive ability can nevertheless get a substantial weight in a combined forecast if the expert’s forecast error is negatively correlated with the forecast errors of the other experts. We therefore relax the independence assumption by also proposing a multi-output GP to jointly model the local predictive ability of a set of $K$ experts over the pooling space.
The multi-output GP uses a kernel function that is constructed by correlating a set of underlying independent GPs, one for each expert. This construction makes it possible to use separate smoothness kernel parameters, or even completely different classes of kernels, for each expert. It also gives us access to interpretable parameters that describe the correlation between experts. The multi-output model allows us to jointly infer the latent GP surfaces of all experts, which gives a more full quantification of the uncertainty of the ELPD surface of prediction pools based on the set of experts.
@@
The joint posterior distribution of the conditional ELPD and kernel hyperparameters are sampled by Hamiltonian Monte Carlo (HMC) using Stan (Stan Development Team,, 2024). The dimension of the sampled conditional ELPD grows with the sample size however, and inferences are time-consuming even for moderately large datasets. We show empirically that a cube-root transformation of the log scores is approximately Gaussian with homoscedastic variance. This makes it possible to integrate out the latent GP surface analytically and sample the low-dimensional marginal posterior kernel hyperparameters by HMC, which is much faster and scalable for larger datasets. We show that the ELPD surface can then be easily sampled by standard GP methods (Rasmussen and Williams,, 2006, Chapter 2) combined with analytically available third moments of Gaussian processes. In the more general case when either the expert or the data generating process is non-Gaussian we propose to use a general power transformation of the ELPD and to estimate the power transformation using HMC jointly with the kernel hyperparameters.
@@
The chapter proceeds as follows. Section 2 defines local predictive ability and describes the noncentral $\chi^{2}$ Gaussian process (GP) regression for estimating it, along with a computationally more efficient single- and multi-output GP regression on power-transformed log scores. Section 3 consists of an empirical example analyzing local predictive ability for a set of models in a bike sharing dataset, along with several methods to pool predictive distributions based on estimated predictive ability. Section 4 concludes.
@@ 2 A framework for estimating local predictive ability
This section defines local predictive ability within a decision maker framework (Oelrich et al.,, 2024), discusses the implications of using log scores as data to estimate local predictive ability, and proposes Gaussian process regression as a model for estimating both joint and marginal local predictive ability.
@@ 2.1 Local predictive ability
The decision maker uses the expected log predictive density for a new data point $\tilde{y}$ from the data-generating process $p_{*}(\cdot)$ to define the predictive ability of expert $k$
@@
where $\tilde{y}$ is a new (out-of sample) observation, $p_{*}(\cdot)$ is the density of the data-generating process, and $\mathbf{y}_{k}$ is the training data the expert has access to. The variable of interest can be univariate or multivariate, but we will for simplicity assume that it is univariate here.
The decision maker further believes that the predictive ability of the experts vary over a set of pooling variables $\mathbf{z}$ (Oelrich et al.,, 2024). She is therefore not only interested in determining how good the experts are at making predictions in general, but how good they are at making predictions conditional on the pooling variables, $\mathbf{z}$. We call this conditional predictive ability the local predictive ability of the expert at $\mathbf{z}$
@@
where the data generating process $p_{*}(\tilde{y}|\mathbf{z})$ may or may not depend on the pooling variables $\mathbf{z}$.
@@ 2.2 Distribution of the log scores
The log scores of the experts, $\ell_{ik}$ for $k=1,\ldots,K$ and $i=1,\ldots,n$, are treated as data by the decision maker (Lindley et al.,, 1979) from which she infers how the ELPD of each expert varies over the pooling space. The choice of pooling variables is therefore a pure modeling problem from the viewpoint of the decision maker. The decision maker does not need to know if the experts have access to the pooling variables when making their predictions, but this information can of course affect the modeling choices of the decision maker.
The distribution of the local predictive ability is determined by the predictive distribution of the expert, which is typically known, and the data-generating process, which is typically unknown. In order to specify a parametric model for the log scores, we therefore need access to the predictive distribution of the expert. Different predictive distribution will lead to different models of local predictive ability.
Consider the practically important scenario where the expert predictive distributions, as well as the data generating process, are all Gaussian. Specifically, let the predictive distribution of expert $k$ for the data point $y_{i}$ be ${y_{i}\mid\mathbf{y}_{k}\sim\mathcal{N}(\mu_{ik},\sigma_{ik}^{2})}$, where $\mathbf{y}_{k}$ is the training data used by the $k$th expert. The log score is then
$\ell_{ik}\equiv\log p_{k}(y_{i}\mid\mathbf{y}_{k})=-\frac{1}{2}\log(2\pi\sigma_{ik}^{2})-\frac{1}{2\sigma_{ik}^{2}}(y_{i}-\mu_{ik})^{2}\;.$ (3)
If we let the data-generating process be given by ${y_{i}\sim\mathcal{N}(\mu_{i*},\sigma_{i*}^{2})}$, we can trivially rewrite Eq. (3) as
$\ell_{ik}\equiv\log p_{k}(y_{i}\mid\mathbf{y}_{k})=-\frac{1}{2}\log(2\pi\sigma_{ik}^{2})-\frac{\sigma_{i*}^{2}}{2\sigma_{ik}^{2}}\left(\frac{y_{i}-\mu_{ik}}{\sigma_{i*}}\right)^{2}.$ (4)
Since $(y_{i}-\mu_{ik})/\sigma_{i*}\sim\mathcal{N}(\mu_{i*}-\mu_{ik},1)$, it follows that (Sivula et al., 2022a, )
@@
where $X_{ik}\sim\chi^{2}_{1}(\lambda_{ik})$ with noncentrality parameter $\lambda_{ik}=(\mu_{i*}-\mu_{ik})^{2}/\sigma_{i*}^{2}$, $a_{ik}=-\frac{1}{2}\log(2\pi\sigma_{ik}^{2})$ and $b_{ik}=\sigma_{i*}^{2}/(2\sigma_{ik}^{2})$. Hence, the noncentrality parameter $\lambda_{ik}$ depends on the difference between the predictive mean of the expert and the mean of the data-generating process; the shape of the distribution of the log scores therefore varies dramatically with $\mu_{i*}-\mu_{ik}$ and is typically highly non-Gaussian for all $b$, see Fig. 1 for an illustration.
@@
where $\chi^{2}_{1}(\lambda,b)$ denotes a scaled noncentral $\chi^{2}$ with one degree of freedom, noncentrality parameter $\lambda$ and scale parameter $b$. A useful simplification of the model in (6) is to assume that $b_{ik}=b_{k}$ for all $i=1,\ldots,n$, i.e. that the variance ratio $\sigma_{i*}^{2}/\sigma_{ik}^{2}$ is constant over observations for a given expert.
2.3 Scaled noncentral $\chi^{2}$ Gaussian process regression
The setup with Gaussian data generating process and experts in the previous subsection suggests a scaled noncentral $\chi^{2}$ regression for modeling the local predictive ability of experts. We will here propose a nonparametric Bayesian model where the log of the noncentrality parameter is modeled by a Gaussian process over the pooling space.
A Gaussian process (GP) generalizes the Gaussian distribution from a random variable to a stochastic process (Rasmussen and Williams,, 2006). In simple terms, a Gaussian process is an infinite sequence of function values with different inputs $\{f(\mathbf{z}_{1}),f(\mathbf{z}_{2}),f(\mathbf{z}_{3}),\dots\}$, any subset of which follows a multivariate Gaussian distribution. Each GP has a covariance function, or kernel, which depends on the input values and specifies how the covariance matrix for each subset of the sequence should be constructed. The specific covariance function we select determines our prior regarding how smoothly the outputs vary over the input space. The covariance function we use in the examples and applications in this chapter is the squared exponential with automatic relevance determination (ARD) (Neal,, 1996)
@@ 2.3 Scaled noncentral $\chi^{2}$ Gaussian process regression
Our noncentral $\chi^{2}$ Gaussian process model for the log scores is
@@
where $\mathcal{N}^{+}(\cdot,\cdot)$ denotes a normal distribution truncated at 00. Note that $b_{k}>0$, and $b_{k}=1/2$ when the variance of the DGP and the expert match. Further, note that the GP is on the log of the noncentrality parameter, ensuring that $\lambda_{k}(\mathbf{z})$ stays non-negative. Since the mean of a $\chi^{2}_{1}(\lambda,b)$ variable is $b(1+\lambda)$, the Gaussian process in (8) directly implies a flexible prior for the mean function of the log scores, i.e. the local ELPD. We will refer to the model in (8) as $\operatorname{GP}(\chi^{2}_{1})$. The mean function $\mu_{k}(\cdot)$ and the covariance kernel $g_{k}(\cdot,\cdot)$ are subscripted by $k$ since their hyperparameters are allowed to be different across the $K$ experts. Any valid kernel function can be used. For notational simplicity, we let $\boldsymbol{\theta}_{k}$ be the vector with all the hyperparameters in the mean and kernel function of the GP prior.
The ultimate goal of the decision maker is to compute the posterior distribution of the local ELPD $\operatorname{\eta}_{k}(\tilde{\mathbf{z}})=\operatorname{E}_{p_{*}}(\tilde{\ell}_{k})=\tilde{a}_{k}-b_{k}(1+\lambda_{k}(\tilde{\mathbf{z}}))$ at a new point $\tilde{\mathbf{z}}$ in pooling space for each expert. This is achieved by sampling from the joint posterior
@@
where $\boldsymbol{\ell}^{\prime}_{k}=(\ell^{\prime}_{1k},\ldots,\ell^{\prime}_{nk})^{\top}$ and $\boldsymbol{\lambda}_{k}=(\lambda_{k}(\mathbf{z}_{1}),\ldots,\lambda_{k}(\mathbf{z}_{n}))^{\top}$ is the vector of noncentrality parameters in the training data. We sample the potentially high-dimensional posterior in (9) using Hamiltonian Monte Carlo (HMC) in Stan (Stan Development Team,, 2024; Gabry and Cesnovar,, 2022).